# by Leveraging CLIP Latent Space

### 07 Feb 2024 v1

@@ Abstract
* indicates equal contribution, $\dagger$ Corresponding Author: maitreya.patel@asu.edu.
Despite the recent advances in personalized text-to-image (P-T2I) generative models, subject-driven T2I remains challenging. The primary bottlenecks include 1) Intensive training resource requirements, 2) Hyper-parameter sensitivity leading to inconsistent outputs, and 3) Balancing the intricacies of novel visual concept and composition alignment. We start by re-iterating the core philosophy of T2I diffusion models to address the above limitations. Predominantly, contemporary subject-driven T2I approaches hinge on Latent Diffusion Models (LDMs), which facilitate T2I mapping through cross-attention layers. While LDMs offer distinct advantages, P-T2I methods’ reliance on the latent space of these diffusion models significantly escalates resource demands, leading to inconsistent results and necessitating numerous iterations for a single desired image. Recently, ECLIPSE has demonstrated a more resource-efficient pathway for training UnCLIP-based T2I models, circumventing the need for diffusion text-to-image priors. Building on this, we introduce $\lambda$-ECLIPSE [^1] The designation $\lambda$-ECLIPSE is inspired by its conceptual alignment with the $\lambda$-calculus. In this context, the $\lambda$-ECLIPSE model functions similarly to a functional abstraction within $\lambda$-calculus, where it effectively binds variables. These variables, in our case, represent novel visual concepts that are integrated through composition prompts.. Our method illustrates that effective P-T2I does not necessarily depend on the latent space of diffusion models. $\lambda$-ECLIPSE  achieves single, multi-subject, and edge-guided T2I personalization with just 34M parameters and is trained on a mere 74 GPU hours using 1.6M image-text interleaved data. Through extensive experiments, we also establish that $\lambda$-ECLIPSE  surpasses existing baselines in composition alignment while preserving concept alignment performance, even with significantly lower resource utilization.
Figure 1: Our $\lambda$-ECLIPSE  can estimate subject-specific image embeddings while maintaining the balance between concept and composition alignment. $\lambda$-ECLIPSE  supports multi-concept personalization and provides additional controls (i.e., edge map) for image generation. In this work, we show that it is possible to perform subject-driven T2I in the latent space of a pre-trained CLIP model without depending on the diffusion UNet models; effectively reducing the heavy resource requirements.
@@ 1 Introduction
The field of text-to-image (T2I) diffusion models has recently witnessed remarkable advancements, achieving greater photorealism and enhanced adherence to textual prompts. This has catalyzed the emergence of diverse applications, notably in controllable T2I models and personalized T2I (P-T2I) models. P-T2I, in particular, encompasses the intricate task of learning and reproducing novel visual concepts or subjects in varied contexts, beyond mere replication of reference images. This personalization task is challenging, necessitating a nuanced balance between concept and compositional alignment. The complexity escalates further when multi-subject personalization is desired.
Early works in this field, such as Textual Inversion [12] and DreamBooth [38], employed per-concept optimization strategies involving fine-tuning certain parameters within T2I diffusion models. However, these methods struggle with generalization challenges and are time-intensive. Contemporary research is pivoting towards fast personalization techniques that involve training hypernetworks and integrating new layers or parameters within pre-trained diffusion UNet models. As summarized in Table 1, despite their relative parameter efficiency, these methods are still resource-intensive (especially for multi-concept personalization), largely due to their reliance on the latent spaces of diffusion models. Additionally, controlling the inherent randomness in diffusion models remains a significant challenge, often resulting in inconsistent outputs and necessitating multiple attempts to generate a single accurate image.
@@
Upon further investigation, we find that most subject-driven T2I approaches build upon variants of the Latent Diffusion Model (LDM) [37], specifically Stable Diffusion models. LDMs employ cross-attention layers to condition diffusion models with text embeddings, necessitating a mapping of target subject images to latent spaces compatible with the diffusion models, either as text embeddings or as supplementary model parameters. This process involves backpropagation through the entire diffusion model, often comprising over a billion parameters, contributing to the inefficiency of existing P-T2I methods.
In contrast, the UnCLIP [35] T2I family of LDMs offers an intriguing alternative, wherein text embeddings are mapped to vision embeddings through a diffusion transformer prior network before passing the conditioning to the diffusion UNet image generator. This mechanism, as illustrated in Figure 2, enables the generator to accurately reproduce original images from their visual embeddings alone. Consequently, our research pivots towards leveraging the UnCLIP T2I stack for P-T2I strategies. However, like their predecessors, these T2I diffusion priors are also parameter-heavy, with around 1 billion parameters. Until recently, ECLIPSE [31] showed that it is possible to compress this 1B parameter of the T2I prior model down to 33M by removing the diffusion modeling.
@@
Table 1: A quick overview of previous works on T2I personalization. Our method is the first to offer multi-concept personalization without depending on diffusion UNet models (except for inference). We provide the extended overview table in the supplementary materials.
Method Multi Finetuning Diffusion Total opt. GPU
Concepts Free Free params Hours
Textual Inversion [12] 768 1
DreamBooth [38] 0.9B 0.5
ELITE [48] 77M -
BLIP-Diffusion [21] 1.5B 2304
IP-Adapter [50] 22M -
@@
Subject-Diffusion [27] 252M -
Kosmos-G [29] 1.9B 12300
$\lambda$-ECLIPSE (ours) 34M 74
Figure 2: CLIP(vision) features capture the semantics and fine-grained visual details. UnCLIP models (e.g., Kandinsky v2.2) can effectively leverage this latent space to reconstruct the image from its embedding.
ECLIPSE posits that text-to-image mapping can be optimized through contrastive pre-training. It inputs text embeddings and estimates corresponding image embeddings, ensuring strong alignment with textual features. Building on these insights, to enhance this framework and deepen the comprehension of novel visual subjects, we introduce a novel pre-training strategy and propose the $\lambda$-ECLIPSE  T2I prior model for the UnCLIP framework. Specifically, we propose a subject-driven instruction tuning task involving the image-text interleaved data. This involves creating 1.6 million high-quality image-text pairs, where text tokens linked to subjects are substituted with image embeddings. Subsequently, $\lambda$-ECLIPSE  is trained to estimate image embeddings that not only harmonize with text semantics but also encapsulate subject representations.
This model facilitates multi-subject-driven image generation using UnCLIP models without imposing excessive resource demands. $\lambda$-ECLIPSE  bypasses the reliance on diffusion latent space and operates within a pre-trained CLIP latent space [33]. We leverage Kandinsky v2.2 diffusion UNet model’s [36] capacity to interpret image structures from CLIP embeddings (refer to Figure 2). Importantly, CLIP image embeddings even accurately capture the details of the canny edge map [6]. Therefore, we extend $\lambda$-ECLIPSE to incorporate canny edge maps as auxiliary guides for subject-driven T2I. Our extensive experimental validation demonstrates that $\lambda$-ECLIPSE  achieves unparalleled performance in single and multi-subject composition alignment while preserving the integrity of reference subjects, competitively rivaling Multimodal Large Language Models (MLLM) based methods that require intense computing resources. At last, we find that $\lambda$-ECLIPSE inherits the continuous latent space from the CLIP; for the first time allowing the interpolations between multiple concepts for P-T2I models. Figure 1 illustrates the qualitative capabilities of $\lambda$-ECLIPSE.
In summary, our contributions are: (1) we introduce a straightforward and yet very effective diffusion-free multi-subject personalization strategy by estimating subject-aligned image embeddings in CLIP(vision) latent space, (2) despite significantly reducing resource requirements, our method achieves state-of-the-art performance, and (3) we demonstrate the potential of utilizing diffusion image generators as standalone pre-trained entities, enabling controlled outputs without necessitating explicit fine-tuning.
2 Related Works
Figure 3: This figure illustrates the three stages of the $\lambda$-ECLIPSE pipeline. 1) Create the image-text interleaved features by utilizing the pre-trained CLIP encoders. 2) Pre-train the $\lambda$-ECLIPSE (34M parameters) using Eq. 1, which ensures the mapping to the desired latent space given the image-text interleaved data. During inference, we just need a single-forward pass to estimate the corresponding image embedding. 3) During inference, the Kandinsky v2.2 diffusion UNet model takes the output from the $\lambda$-ECLIPSE (prior) and generates the corresponding images.
@@ 2 Related Works
Pioneering efforts in image generation, notably DALL-E [34] and CogView [11], leveraged autoregressive models to achieve significant results. Recent advancements predominantly feature diffusion models, acclaimed for their high image fidelity and diversity in text-to-image (T2I) generation. A notable example is Stable Diffusion, which builds upon the Latent Diffusion Model (LDM) [37] and excels in semantic and conceptual understanding by transitioning training to latent space. Imagen [40], Pixart-$\alpha$ [8], and DALL-E 3 [5] propose using a large T5 language model to improve language understanding. DALL-E 2 [35] along with its UnCLIP variation models such as Kandinsky [36] and Karlo, uses a diffusion prior and diffusion UNet modules to generate images using the pre-trained CLIP [33] model.
@@
Approaches like Textual Inversion [12], DreamBooth [38], and Custom Diffusion [20] focus on training specific parameters to encapsulate visual concepts. LoRA [16] and Perfusion [45] target efficient fine-tuning adjustments, particularly rank 1 modifications. However, these methods are constrained by their requirement for concept-specific tuning. ELITE [48] was the first approach addressing fast personalization for single-subject T2I. BLIP-Diffusion [21] adapts the BLIP2 encoder [22], training approximately 1.5B parameters to enable zero-shot, subject-driven image generation. IP-Adapter [50] introduces a decoupled cross-attention mechanism, negating the need to train the foundational UNet model by permitting fine-tuning of a reduced parameter set (22M).
Mix-of-Show [14] and Zip-LoRA [42] train individual concepts and then combine them to generate multiple subjects. Break-A-Scene [4] shows multi-concept capability but requires single images containing diverse objects. Subject Diffusion [27] creates a high-quality dataset and presents the precision control for fast multi-subject image generation. Kosmos-G [29], akin to Subject-Diffusion, employs a Multimodal Large Language Model (MLLM) for text-image embedding alignment, though it necessitates extensive parameter optimization (1.9B). These multi-subject P-T2I methods are not only demanding in terms of parameters but also depend on a massive number of frozen parameters of the diffusion UNet model, increasing training computational loads. In contrast, our model, $\lambda$-ECLIPSE, forgoes test-time fine-tuning and training-time reliance on the diffusion UNet model for single and multi-concept, control-guided P-T2I, positioning it as a resource-efficient solution.
At last, methods like GLIGEN [23], ControlNet [51], and UniControl [32] incorporate additional controls (i.e., edge map, depth, segmentations) into the diffusion model to generate the desired images. BLIP-Diffusion, IP-Adapter, and Kosmos-G can leverage such pre-trained controls. However, in many scenarios, these controls are too strong which makes generated images lose subject-specific details. While $\lambda$-ECLIPSE can be plugged with these pre-trained methods, we explore the possibility of learning canny edge as the additional guidance for P-T2I. We show that $\lambda$-ECLIPSE learns to balance the edge map, subjects, and text composition.
@@ 3 Method
In this section, we introduce $\lambda$-ECLIPSE, our approach to multi-subject text-to-image personalization. This method extends the capabilities of the existing ECLIPSE  framework through a novel image-text interleaved pretraining strategy, notably omitting the need for explicit diffusion modeling. Our approach mainly capitalizes on the efficient utilization of the CLIP latent space. Figure 3 outlines the end-to-end framework.
The primary objective of $\lambda$-ECLIPSE is to facilitate both single and multi-subject T2I generation processes, accommodating controlled conditional guidance such as edge maps. Initially, we detail the problem formulation and elaborate on the extended design of the ECLIPSE  T2I prior module. Subsequently, we delve into the image-text interleaved training methodology. This fine-tuning process enables the ECLIPSE  model to harness semantic correlations between CLIP image and text latent spaces while preserving subject-specific visual features. Upon attaining the target embedding, the diffusion UNet model, derived from Kandinsky v2.2, is employed to reconstruct the image.
@@ 3.1 Text-to-Image Prior Mapping
In the UnCLIP T2I models, the objective of the text-to-image prior model ($f_{\theta}$) is to establish a proficient text-to-image embedding mapping. This model is designed to adeptly map textual representations to their corresponding visual embeddings, denoted as ($f_{\theta}:z_{y}\rightarrow z_{x}$), where $z_{x/y}$ represent the embeddings for images and text, respectively. The visual embedding predictions ($\hat{z_{x}}=f_{\theta}(z_{y})$) are then effectively utilized by the diffusion image generators ($h_{\phi}$), which are inherently conditioned on these vision embeddings ($h_{\phi}:z_{x}\rightarrow x$).
As illustrated in Figure 2, these diffusion image generators possess a remarkable capability for precise image reproduction, since $z_{x}$ encompasses comprehensive image information. Our goal is to accurately estimate the image embedding $\hat{z}_{x}$, incorporating the subject representations, thereby eliminating reliance on $h_{\phi}$ during training. Existing LDM-based P-T2I methods are limited by the LDM’s singular module approach ($h_{\phi}:z_{y}\rightarrow x$). Consequently, mastering the latent space of $h_{\phi}$ becomes essential for effective P-T2I for the baseline methodologies.
We propose a mapping function, $f^{\prime}_{\theta}$, which processes text representations ($z_{y}$) alongside subject ($x_{i}$) specific representations ($z_{x_{i}}$), to derive an image embedding that encapsulates both text prompts and subject visuals ($\hat{z}_{x}$). The challenge lies in harmonizing $z_{x_{i}}$ and $z_{y}$ within $f^{\prime}_{\theta}:(z_{y},z_{x_{i}})\rightarrow\hat{z}_{x}$, ensuring alignment while preventing overemphasis on either aspect, as this could compromise composition alignment. To address this, we employ the contrastive pre-training strategy after [31]:
$\begin{split}\mathcal{L}_{prior}=&\underset{\begin{subarray}{c}\epsilon\sim\mathcal{N}(0,I)\\ z_{y},z_{x}\end{subarray}}{\operatorname{\mathbb{E}}}\Big{[}||z_{x}-f_{\theta}(\epsilon,z_{y})||_{2}^{2}\Big{]}\\ &-\frac{\lambda}{N}\sum_{i=0}^{N}\log\frac{\exp(\langle\hat{z}^{i}_{x},z^{i}_{y}\rangle/\tau)}{\sum_{j\in[N]}\exp(\langle\hat{z}^{i}_{x},z^{j}_{y}\rangle/\tau)}.\end{split}$ (1)
Here, $\lambda$ serves as the hyperparameter. The first loss term measures the mean-squared error between the estimated and actual image embeddings, primarily ensuring concept alignment. However, our preliminary studies reveal that exclusive reliance on this term diminishes composition alignment, echoing observations made by [31]. Therefore, we stick with the contrastive loss component to bolster compositional generalization, with $\lambda$ balancing concept and composition alignment.
Figure 4: Qualitative examples of $\lambda$-ECLIPSE.
Figure 5: This figure illustrates the qualitative comparison of $\lambda$-ECLIPSE w.r.t. the state-of-the-art approaches for single-subject T2I generations. We choose concepts and prompts from the Dreambench dataset. For all methods, concepts, and prompts, we generate the four images and select the best one that represents the concept and composition more accurately.
@@
Acknowledging the limitations in existing methods, which necessitate learning the diffusion latent space even for additional control inputs, we endeavor to achieve a more nuanced balance between subject, text, and supplementary conditions. Consequently, we have augmented $\lambda$-ECLIPSE to accommodate an additional modality, a canny edge map, providing more refined control over subject-driven image generation. This entails modifying the prior model to accept additional conditions ($f_{\theta}^{\prime}:(z_{y},z_{x_{i}},z_{c})\rightarrow\hat{z}_{x}$, where $z_{c}$ symbolizes the additional modality embedding). During training, we drop $z_{c}$ for 1% to improve the unconditional generations.
Incorporating canny edge maps during training enhances stability and broadens the generalization capabilities of $\lambda$-ECLIPSE, yielding benefits even in the absence of these controls during inference. Our findings indicate that $\lambda$-ECLIPSE successfully learns a unified mapping function, effectively estimating target image representations from a synergy of text, subjects, and edge map.

### 09 Apr 2024 v2

# Maitreya $*{}^{\dagger}$   Sangmin Jung*   Chitta Baral   Yezhou Yang

@@ Abstract
Despite the recent advances in personalized text-to-image (P-T2I) generative models, it remains challenging to perform finetuning-free multi-subject-driven T2I in a resource-efficient manner. Predominantly, contemporary approaches, involving the training of Hypernetworks and Multimodal Large Language Models (MLLMs), require heavy computing resources that range from 600 to 12300 GPU hours of training. These subject-driven T2I methods hinge on Latent Diffusion Models (LDMs), which facilitate T2I mapping through cross-attention layers. While LDMs offer distinct advantages, P-T2I methods’ reliance on the latent space of these diffusion models significantly escalates resource demands, leading to inconsistent results and necessitating numerous iterations for a single desired image. In this paper, we present $\lambda$-ECLIPSE, an alternative prior-training strategy that works in the latent space of a pre-trained CLIP model without relying on the diffusion UNet models. $\lambda$-ECLIPSE leverages the image-text interleaved pre-training for fast and effective multi-subject-driven P-T2I. Through extensive experiments, we establish that $\lambda$-ECLIPSE surpasses existing baselines in composition alignment while preserving concept alignment performance, even with significantly lower resource utilization. $\lambda$-ECLIPSE performs multi-subject driven P-T2I with just 34M parameters and is trained on a mere 74 GPU hours. Additionally, $\lambda$-ECLIPSE demonstrates the unique ability to perform multi-concept interpolations.
* indicates equal contribution, $\dagger$ corresponding author: maitreya.patel@asu.edu
Figure 1: $\lambda$-ECLIPSE  can estimate subject-specific image embeddings while maintaining the balance between concept and composition alignment in a resource-efficient way.
@@ 1 Introduction
The field of text-to-image (T2I) diffusion models has recently witnessed remarkable advancements, achieving greater photorealism and enhanced adherence to textual prompts. This has catalyzed the emergence of diverse applications, notably subject-driven personalized T2I (P-T2I) models. In particular, this encompasses the intricate task of learning and reproducing novel visual concepts or subjects in varied contexts requiring high concept and compositional alignment. The complexity escalates further when multi-subject personalization is desired.
Early works employed concept-specific optimization strategies involving fine-tuning certain parameters within T2I diffusion models gal2022image ; ruiz2023dreambooth ; kumari2023multi ; tumanyan2023plug ; gal2023encoder . Although these methods achieve state-of-the-art (SOTA) performance, they struggle with generalization and are time-intensive. Contemporary research is pivoting towards fast personalization techniques. Within this paradigm, there are two types of popular approaches: 1) Methods that involve training hypernetworks and integrating new layers or parameters within pre-trained diffusion UNet models Wei2023ELITEEV ; ye2023ip ; tewel2023key ; shah2023ziplora ; ruiz2023hyperdreambooth , and 2) MLLM-based learning of prior models that focuses on leveraging text-latent space of frozen diffusion UNet model pan2023kosmos ; sun2023generative .
The hypernetwork-based strategy achieves single-concept customization but has not been extended to multi-concepts. Moreover, when combined with additional control (i.e. Canny edge map), they struggle to maintain the concept alignment ($\sim$30drop in performance; Section 4) and strongly favor the edge map. At the same time, MLLM-based approaches can perform fast multi-concept customization but require heavy computing resources. In Table 1, we provide the overview of various single and multi-concept customization methodologies in terms of total parameters, iterations, and GPU hours required to train the models. It can be observed that multi-concept customization methodologies further increase the resource requirements. For instance, Kosmos-G pan2023kosmos consumes 18x more resources than IP-Adapter ye2023ip . And Emu2 sun2023generative requires training of 19x more parameters compared to Kosmos-G. Hence, despite MLLMs’ seemingly useful scenarios it is not viable to blindly train them. Therefore, in this work, we focus on answering one question: What are the alternative methodology and design choices we can make to improve resource efficiency?
@@
Upon further investigation, we find that most subject-driven T2I approaches build upon variants of the Latent Diffusion Model (LDM) rombach2022high , specifically Stable Diffusion models. These LDMs employ cross-attention layers to condition diffusion models with text embeddings, necessitating a mapping of target subject images to latent spaces compatible with the diffusion models at the prior training stage. This is also known as score distillation instruction tuning for MLLMs pan2023kosmos . As there is no choice but to learn this text-to-image diffusion latent space, it involves backpropagation through the entire diffusion model often comprising over a billion parameters, contributing to the inefficiency of existing P-T2I methods.
To improve the resource efficiency for multi-concept image generation, we present $\lambda$-ECLIPSE [^1] The designation $\lambda$-ECLIPSE is inspired by its conceptual alignment with the $\lambda$-calculus. In this context, the $\lambda$-ECLIPSE model functions similarly to a functional abstraction within $\lambda$-calculus, where it effectively binds variables. These variables, in our case, represent novel visual concepts that are integrated through composition prompts. Here, ECLIPSE indicates our architecture design choice., which leverages the properties of UnCLIP T2I models (e.g. DALL-E 2 ramesh2022hierarchical and Kandinsky v2.2 razzhigaev2023kandinsky ) and performs P-T2I in the compressed latent space. Specifically, unlike previous MLLM-based methodologies, $\lambda$-ECLIPSE aligns the output space of priors with CLIP vision space instead of the CLIP text space. $\lambda$-ECLIPSE takes multiple images and text instructions as input and estimates the respective vision embeddings, which can be used by the frozen diffusion UNet model from the UnCLIP stack to generate the resulting image. This elevates the training time dependencies on diffusion models for P-T2I; significantly contributing to the resource efficiency. Additionally, as diffusion or MLLM-based priors are still compute heavy due to a huge number of parameters and slower convergence, we build upon ECLIPSE patel2023eclipse and SEED ge2023making , which shows that text-to-image mapping can be optimized through contrastive pre-training. Here, we select ECLIPSE as preferred choice of prior architecture for best efficiency. At last, we propose a subject-driven instruction tuning task based on the image-text interleaved data as a pre-training strategy. This involves creating 2 million high-quality image-text pairs, where text embeddings linked to subjects are substituted with the respective image embeddings, which in return are considered as input to the $\lambda$-ECLIPSE. While $\lambda$-ECLIPSE can be plugged with these pre-trained methods, we explore the possibility of $\lambda$-ECLIPSE to incorporate Canny edge as an additional control to synergetically work with subject-driven image generation tasks. Figure 1 provides the overview of $\lambda$-ECLIPSE capabilities.
Overall, we propose $\lambda$-ECLIPSE as an initial attempt to motivate future works on designing resource-efficient solutions for MLLM-based approaches. We summarize our main contributions as follows: 1) We introduce a training-time diffusion-independent, UnCLIP-based prior learning strategy for enabling efficient and fast multi-subject customization. 2) Extensive experiments on Dreambench, Multibench, and ConceptBed reveal that $\lambda$-ECLIPSE (34M parameter model) trained on a mere 74 GPU hours can achieve competitive performance to that of big counterparts (having 2B-37B parameters) and improve text-composition alignment. 3) At last, $\lambda$-ECLIPSE inherits the smooth CLIP latent space. This allows us to perform the seamless transition between multi-concept generated images.
2 Related Works
@@
Table 1: A quick overview of previous works on P-T2I. Our method is the first to offer multi-concept-driven generation without depending on diffusion UNet models (except for inference). We provide the extended overview table in the appendix.
Method Multi Finetuning Diffusion Total opt. Training Dataset GPU
Concepts Free Free params Steps Size Hours
Textual Inversion gal2022image 768 5000 - 1
DreamBooth ruiz2023dreambooth 0.9B 800 - 0.2
ELITE Wei2023ELITEEV 77M 135K 125K 336
BLIP-Diffusion li2023blip 1.5B 500K 129M 2304
IP-Adapter ye2023ip 22M 1M - 672
@@
Subject-Diffusion ma2023subject 252M 300K 76M -
Kosmos-G pan2023kosmos 1.9B 800K 200M 12300
Emu-2 sun2023generative 37B 70K 162M -
$\lambda$-ECLIPSE (ours) 34M 100K 2M 74
@@
Pioneering efforts in image generation, notably DALL-E ramesh2021zero and CogView ding2021cogview , leveraged autoregressive models to achieve significant results. Recent advancements predominantly feature diffusion models, acclaimed for their high image fidelity and diversity in text-to-image (T2I) generation. A notable example is Stable Diffusion, which builds upon the Latent Diffusion Model (LDM) rombach2022high and excels in semantic and conceptual understanding by transitioning training to latent space. Imagen saharia2022photorealistic , Pixart-$\alpha$ chen2023pixart , and DALL-E 3 betker2023improving propose using a large T5 language model to improve language understanding. DALL-E 2 ramesh2022hierarchical along with its UnCLIP variation models such as Kandinsky razzhigaev2023kandinsky and Karlo kakaobrain2022karlo-v1-alpha , uses a diffusion prior and diffusion UNet modules to generate images using the pre-trained CLIP radford2021learning model.
@@
Approaches like Textual Inversion gal2022image , DreamBooth ruiz2023dreambooth , and Custom Diffusion kumari2023multi focus on training specific parameters to encapsulate visual concepts. LoRA hu2021lora and Perfusion tewel2023key target efficient fine-tuning adjustments, particularly rank 1 modifications. However, these methods are constrained by their requirement for concept-specific tuning. ELITE Wei2023ELITEEV was the first approach addressing fast customized generation for single-subject T2I. BLIP-Diffusion li2023blip adapts the BLIP2 encoder li2023blip2 , training approximately 1.5B parameters to enable zero-shot, subject-driven image generation. IP-Adapter introduces a decoupled cross-attention mechanism, negating the need to train the foundational UNet model by permitting fine-tuning of a reduced number of 22M parameters.
Mix-of-Show gu2023mix and Zip-LoRA shah2023ziplora train individual concepts and then combine them to generate multiple subjects. Break-A-Scene avrahami2023break shows multi-concept capability but requires single images containing diverse objects. Subject Diffusion ma2023subject creates a high-quality dataset and presents the precision control for fast personalized multi-subject image generation. Kosmos-G and Emu2 sun2023generative , akin to Subject-Diffusion ma2023subject , employs a Multimodal Large Language Model (MLLM) for text-image embedding alignment, though it necessitates extensive parameter optimization (1.9B-37B). These multi-subject P-T2I methods are not only demanding in terms of parameters but also depend on a massive number of frozen parameters of the diffusion UNet model, increasing training computational loads. In contrast, our model, $\lambda$-ECLIPSE, forgoes test-time fine-tuning and training-time reliance on the diffusion UNet model for single and multi-concept, control-guided P-T2I, positioning it as a resource-efficient solution.
At last, methods like GLIGEN li2023gligen , ControlNet zhang2023adding , and UniControl qin2023unicontrol incorporate additional controls (i.e., edge map, depth, segmentations) into the diffusion model to generate the desired images. BLIP-Diffusion, IP-Adapter, and Kosmos-G can leverage such pre-trained controls. However, in many scenarios, these controls are too strong, making generated images lose subject-specific details. We show that $\lambda$-ECLIPSE learns to balance the edge map, subjects, and composition. We offer a more comprehensive review of related works in the appendix.
@@ 3 Method
Figure 2: This figure illustrates the three stages of the $\lambda$-ECLIPSE pipeline. 1) Create the image-text interleaved features using frozen CLIP. 2) Pre-train the $\lambda$-ECLIPSE (34M parameters) using Eq. 1, which ensures the mapping to the desired latent space given the image-text interleaved data. 3) During inference, the frozen Kandinsky v2.2 diffusion UNet model takes the output from the $\lambda$-ECLIPSE  and generates the image.
In this section, we introduce $\lambda$-ECLIPSE, our approach to multi-subject personalized text-to-image generation. Our method combines the contrastive text-to-image strategy from ECLIPSE with the novel image-text interleaved pretraining strategy, notably omitting the need for explicit diffusion modeling. Our approach mainly capitalizes on the efficient utilization of the CLIP latent space. Figure 2 outlines the end-to-end framework.
The primary objective of $\lambda$-ECLIPSE is to facilitate single and multi-subject P-T2I generation processes, accommodating edge maps as conditional guidance. Initially, we detail the problem formulation and elaborate on the UnCLIP stack design of the $\lambda$-ECLIPSE. Subsequently, we delve into the image-text interleaved training methodology. This fine-tuning process enables the $\lambda$-ECLIPSE to harness semantic correlations between CLIP image and text latent spaces while preserving subject-specific visual features.
@@ 3.1 Text-to-Image Prior Mapping
Figure 3: CLIP(vision) features capture the semantics and fine-grained visual details. Each input is given as input to the Kandinsky v2.2 and re-generated from the decoder. (Top: Real-images, Bottom: Canny edge)
In the UnCLIP T2I models, the objective of the text-to-image prior model ($f_{\theta}$) is to establish a proficient text-to-image embedding mapping. This model is designed to adeptly map textual representations to their corresponding visual embeddings, denoted as ($f_{\theta}:z_{y}\rightarrow z_{x}$), where $z_{x/y}$ represent the embeddings for images and text, respectively. The visual embedding predictions ($\hat{z_{x}}=f_{\theta}(z_{y})$) are then effectively utilized by the diffusion image generators ($h_{\phi}$), which are inherently conditioned on these vision embeddings ($h_{\phi}:z_{x}\rightarrow x$). In our experiments, we utilize the Kandinsky v2.2 diffusion UNet model as $h_{\phi}$.
As shown in Figure 3, the CLIP vision encoder preserves the necessary features in $z_{x}$ required to reconstruct the input image and achieves a similar concept alignment score (DINO: 0.66) as finetuning-based DreamBooth method ruiz2023dreambooth .
Our goal is to accurately estimate the image embedding $\hat{z}_{x}$, incorporating the subject representations, thereby eliminating reliance on $h_{\phi}$ during training. Existing LDM-based P-T2I methods are limited by the LDM’s singular module approach ($h_{\phi}:z_{y}\rightarrow x$). Consequently, mastering the latent space of $h_{\phi}$ becomes essential for effective P-T2I for the baseline methodologies, which limits the previous methodologies.
We propose a new mapping function, $f_{\theta}$, which processes text representations ($z_{y}$) alongside subject ($x_{k}$) specific visual representations ($z_{x_{k}}$), to derive an image embedding that encapsulates both text prompts and subject visuals ($\hat{z}_{x}$). The challenge lies in harmonizing $z_{x_{k}}$ and $z_{y}$ within $f_{\theta}:(z_{y},z_{x_{k}})\rightarrow\hat{z}_{x}$, ensuring alignment while preventing overemphasis on either aspect, as this could compromise composition alignment. To address this, we employ the contrastive pre-training strategy after patel2023eclipse :
$\begin{split}\mathcal{L}_{prior}=\underset{\begin{subarray}{c}\epsilon\sim\mathcal{N}(0,I)\\ z_{y},z_{x_{k}}\end{subarray}}{\mathbb{E}}\Big{[}||z_{x}-f_{\theta}(\epsilon,z_{y},z_{x_{k}})||_{2}^{2}\Big{]}-\frac{\lambda}{N}\sum_{i=0}^{N}\log\frac{\exp(\langle\hat{z}^{i}_{x},z^{i}_{y}\rangle/\tau)}{\sum_{j\in[N]}\exp(\langle\hat{z}^{i}_{x},z^{j}_{y}\rangle/\tau)}.\end{split}$ (1)
[…]
@@
Acknowledging the limitations in existing methods, which necessitate learning the diffusion latent space even for additional control inputs, we endeavor to achieve a more nuanced balance between subject, text, and supplementary conditions. Consequently, we have augmented $\lambda$-ECLIPSE to accommodate an additional modality, a Canny edge map, providing more refined control over subject-driven image generation. This entails modifying the prior model to accept additional conditions ($f_{\theta}^{\prime}:(z_{y},z_{x_{i}},z_{c})\rightarrow\hat{z}_{x}$, here $z_{c}$ symbolizes the additional modality embedding).
During training, we drop $z_{c}$ for 1% to improve the unconditional generations. This enhances the stability and broadens the generalization capabilities of $\lambda$-ECLIPSE, yielding benefits even in the absence of these controls during inference. Our results demonstrate that $\lambda$-ECLIPSE  learns a unified mapping function, accurately estimating target image representations through the effective integration of text, image, and edge maps.